# Homogenization with Guaranteed Bounds via Primal-Dual Physically Informed Neural Networks

### v1

@@ 1 Introduction
Accurately modelling the effective behaviour of materials with complex microstructures is crucial in various scientific and engineering domains. Simulation of components made of such materials, while resolving all microstructural details, makes multi-scale modeling prohibitively computationally demanding for practical everyday use. Homogenization aims to replace a heterogeneous material with a homogeneous one that exhibits an equivalent macroscopic behaviour (Engquist and Souganidis, 2008). For most computational homogenization problems, physics-informed neural networks (PINNs) (Raissi et al., 2019) currently cannot match the computational efficiency and reliability of classical Finite Element Method (FEM) and Fast Fourier Transform (FFT) approaches (Geers et al., 2010; Davoli et al., 2024), despite promising theoretical advantages (Grossmann et al., 2024a). However, PINNs demonstrate clear advantages for mesh-free computation of complex microstructures, multi-query parametric studies, and scenarios involving irregular geometries where traditional meshing becomes problematic (Rodrigues et al., 2018).
@@
Recent advances in PINN architectures have begun addressing fundamental limitations in multiscale problems (Park and Zhu, 2022; Leung et al., 2022; Soyarslan and Pradas, 2024). These developments, combined with the inherent advantages of mesh-free computation and rapid evaluation for trained networks, position PINNs as valuable tools for specific applications while classical methods remain optimal for computational homogenization workflows in production.
@@
A critical gap remains in establishing rigorous error bounds and robust formulations for elliptic partial differential equations (PDEs) within the PINN framework. While existing PINN approaches primarily rely on the strong form of the PDE residual, this formulation can be sensitive to solution regularity and may not provide reliable error estimates (Krishnapriyan et al., 2021; Ryck et al., 2023; Mishra and Molinaro, 2023). Variational PINNs (VPINNs) address some of these limitations by incorporating weak formulations that utilize test functions and integrate the PDE residual over the domain, thereby imposing weaker regularity requirements and potentially enhancing solution stability. However, VPINNs remain sensitive to the choice of test functions and still lack robustness in error estimation (E and Yu, 2018; Kharazmi et al., 2019).
@@
To address these challenges in PINN-based homogenization, we propose a scheme that enables the calculation of guaranteed upper and lower bounds for effective parameters. Using a simple microstructure example, we demonstrate how classic PINNs can fail when applied to piecewise constant materials. Although sophisticated techniques exist for handling discontinuous materials, including GFEM (Legrain et al., 2011), X-FEM (Belytschko et al., 2009), the immersed boundary method (LeVeque and Li, 1994), domain decomposition PINNs (Ameya D. Jagtap and George Em Karniadakis, 2020), and various adaptive PINN approaches (Liu et al., 2023; Zhao and Shao, 2025), robust application of advanced PINNs in homogenization settings remains under development.
@@
Our approach employs both primal and dual formulations of the elliptic PDE governing the homogenization problem, rather than relying solely on the primal formulation. The dual formulation provides complementary information that enables more accurate and reliable bounds on effective material properties, as demonstrated for classical homogenization methods by Gaynutdinova et al. (2022, 2023). For the network architecture, we adopt the frameworks of Jiang et al. (2023) and Soyarslan and Pradas (2024), which incorporate periodic boundary conditions directly without loss penalty terms. We extend their architecture to solve the dual formulation problem while satisfying the divergence-free constraint (Richter-Powell et al., 2022; Farkane et al., 2023), then use the primal and dual solutions to derive error bounds. Recognizing that the strong form can still yield poor performance in classic PINNs, we also implement the primal-dual approach for VPINNs and compare results across different test function choices (Berrone et al., 2022; Rojas et al., 2024). Our results reveal that strong-form PINNs outperform both VPINNs and FEM benchmarks only when materials exhibit sufficiently gradual phase transitions, while VPINNs show sensitivity to test function selection. In all cases, however, the guaranteed bounds to the effective parameters can be utilized to assess the quality of the computed results.
@@
The rest of this work is structured as follows. In Section 2, we recall the zeroth-order homogenization theory for both primal and dual formulations in the steady-state linear 2D heat equation. In Section 3, we introduce the PINN architecture and construct the strong training loss for the dual formulation, as well as the variational loss for both primal and dual formulations. In Subsection 4.1, we compare the performance of the classic PINNs for different material functions and address a common mode of PINN failure. Subsection 4.2 compares the performance of VPINNs for different types and cardinalities of test bases. Section 5 summarizes our findings and discusses the pros and cons of each method, and Section 6 provides the outlook for future applications.
@@ 2 Primal and Dual Formulation of Homogenization
In this part, we employ computational homogenization, which determines the effective properties of heterogeneous materials by solving boundary value problems on their representative periodic cells. First, we introduce the governing equation for the homogenization of linear steady-state heat transport in 2D by recalling the strong form of the equation, from which the typical PINN loss is constructed (Park and Zhu, 2022; Leung et al., 2022; Jiang et al., 2023; Wu et al., 2023; Soyarslan and Pradas, 2024), and introduce its dual form. Then we present the variational forms for both the primal and dual formulations, which are used in most classical numerical methods, and will serve as the basis for the VPINN approach. Finally, we show how to calculate the estimates of the homogenized conductivity matrix, as well as its guaranteed upper and lower bounds.
@@ 2.1 Strong form
Let us consider a periodic unit cell $X\subset\mathbb{R}^{2}$ representing a periodic heterogeneous solid, where its microstructural thermal conductivity second-order tensor is defined by a symmetric matrix function $\bm{A}(\bm{x}):X\to\mathbb{R}_{\rm{sym}}^{2\times 2}$, which is uniformly positive definite, with each component essentially bounded in $X$. According to the zeroth-order homogenization theory (Michel et al., 1999; Jikov et al., 2012), the temperature field $u(\bm{x})$ within the unit cell takes the form
$u(\bm{x})=\bm{\xi}^{\rm{T}}\bm{x}+\widetilde{u}(\bm{x}),\;\bm{x}\in X,$
where $\bm{\xi}^{\rm{T}}\bm{x}$ represents the averaged contribution for the prescribed macroscopic temperature gradient $\bm{\xi}$ and $\widetilde{u}(\bm{x})$ is the micro-fluctuation part caused by the material heterogeneity. The temperature fluctuation $\widetilde{u}(\bm{x})$ is a periodic function in the $x_{1}$$x_{2}$ plane, $\bm{x}=[x_{1},x_{2}]$.
Under steady-state conditions, the heat flux $\bm{j}(\bm{x})$ is related to the temperature gradient $\bm{e}(\bm{x})=\nabla u(\bm{x})=[\frac{\partial u}{\partial x_{1}},\frac{\partial u}{\partial x_{2}}]^{\rm T}$ by Fourier’s law as:
$\bm{j}(\bm{x})=-\bm{A}(\bm{x})\bm{e}(\bm{x})=-\bm{A}(\bm{x})\nabla u(\bm{x}).$
The strong form of the governing differential equation of the steady-state heat conduction then requires one to find $\widetilde{u}\in C^{2}_{\rm per}(X)$ (twice differentiable periodic continuous function on $X$) such that:
@@
where $\nabla\cdot$ denotes the divergence operator, $\nabla\cdot\bm{f}=\frac{\partial f_{1}}{\partial x_{1}}+\frac{\partial f_{2}}{\partial x_{2}}$.
The dual form of Eq. (1) is defined for the thermal resistivity $\bm{A}^{-1}(\bm{x})$ (Vondřejc et al., 2015), and its counterpart in electrostatics is commonly known from Maxwell’s equations (Griffiths, 2017). For the heat transfer problem, we can derive it analogously from the property of the temperature gradient field $\bm{e}$ (Briane and Manceau, 2008; Milton, 2022), following the Helmholtz decomposition theorem (Jikov et al., 2012), i.e.:
$\nabla\times\bm{e}(\bm{x})=\bm{0},\;\forall\bm{x}\in X.$
@@
$\widetilde{\bm{v}}(\bm{x})=\bm{Q}\nabla\widetilde{w}(\bm{x}),$
where $\bm{Q}$ is a $90^{\circ}$ rotation matrix on the $x_{1}$$x_{2}$ plane (Girault and Raviart, 1979; Milton, 2022). Then problem (2) can be written as to find $\widetilde{w}(\bm{x})\in C^{2}_{\rm per}(X)$ such that
@@
and the divergence-free constraint (3) is fulfilled automatically. Note that when the curl operator $\nabla\times$ operates on a 2D vector function $\bm{f}$, it is expanded into a 3D vector function $f_{1},f_{2},0]$. The condition $\nabla\times\bm{f}=\bm{0}$ then can be written as
$\left[\frac{\partial f_{3}}{\partial x_{2}}-\frac{\partial f_{2}}{\partial x_{3}},\frac{\partial f_{1}}{\partial x_{3}}-\frac{\partial f_{3}}{\partial x_{1}},\frac{\partial f_{2}}{\partial x_{1}}-\frac{\partial f_{1}}{\partial x_{2}}\right]=\bm{0}.$
Since $f_{3}=0$ and $\bm{f}$ is a function on the $x_{1}$$x_{2}$ plane, the first two components are always zero, and the equation (4) simplifies to contain only the last component, thus reducing to a scalar problem. We keep the original notation for brevity, but from now on, when we write $\nabla\times\bm{f}$, we actually mean its scalar form $\frac{\partial f_{2}}{\partial x_{1}}-\frac{\partial f_{1}}{\partial x_{2}}$.
@@ 2.2 Weak form
Despite promising theoretical advantages, strong-form PINNs currently cannot match the computational efficiency and reliability of classical computational approaches for most PDEs (Grossmann et al., 2024b). As the analytical solution of problems (1) and (4) is usually unavailable, most frameworks, such as the finite element method (FEM) (Yvonnet, 2019) and Fourier-Galerkin method (Moulinec and Suquet, 1994; Schneider, 2021), rely on their respective variational (weak) forms. This allows for the relaxation of the solution regularity requirement from $C^{2}_{\rm per}(X)$ to $H^{1}_{\rm per}(X)$ space. Instead of solving (1), we want to find $\widetilde{u}\in H^{1}_{\rm per}(X)$ such that
$\int_{Y}(\nabla\phi(\bm{x}))^{\rm{T}}\bm{A}(\bm{x})(\bm{\xi}+\nabla\widetilde{u}(\bm{x})){\rm d}\bm{x}=0,\;\;\text{for all }\phi\in H^{1}_{\rm per}(X),$ (5)
where $\phi(\bm{x})$ is a test function. Similarly, the variational form of (4) is to find $\widetilde{w}\in H^{1}_{\rm per}(X)$ such that
$\int_{Y}(\bm{Q}\nabla\psi(\bm{x}))^{\rm{T}}\;\bm{A}^{-1}(\bm{x})(\bm{\xi}+\bm{Q}\nabla\widetilde{w}(\bm{x})){\rm d}\bm{x}=0,\;\;\text{for all }\psi\in H^{1}_{\rm per}(X),$ (6)
with a test function $\psi(\bm{x})$.
@@ 2.3 Upper and lower bound to homogenized parameters
The exact solution of (5) is also the minimum of the related homogenization problem (Jikov et al., 2012), which reads to find a homogenized coefficient matrix $\bm{A}^{*}\in{\mathbb{R}}^{2\times 2}$ such that
@@
Conversely, the inverse $\bm{B}^{*}$ of the homogenized coefficient $\bm{A}^{*}$ can be calculated using the dual formulation (Haslinger and Dvořák, 1995; Vondřejc et al., 2015; Gaynutdinova et al., 2022):
@@
Given the variational structure of (7) and (8), substituting any approximate solutions $\widetilde{u}_{h}$ and $\widetilde{w}_{h}$ into the integrals in (7) and (8), respectively, for given $\bm{\xi}$ and $\bm{\zeta}$ yields larger values than for the respective minimizers, and thus the upper bounds $\bm{A}^{*}_{h}$ and $\bm{B}^{*}_{h}$ to $\bm{A}^{*}$ and $\bm{B}^{*}$, in the sense that
@@
In practice, we solve (1) or (5) (as well as (4) and (6)) for $\bm{\xi},\bm{\zeta}\in\{[1,0]^{\rm T},[0,1]^{\rm T}\}$ to provide the diagonal elements of $\bm{A}^{*}_{h}$ and $\bm{B}^{*}_{h}$.

### v2

@@ 1 Introduction
Accurately modelling the effective behaviour of materials with complex microstructures is crucial in various scientific and engineering domains. Simulation of components made of such materials, while resolving all microstructural details, makes multi-scale modeling prohibitively computationally demanding for practical everyday use. Homogenization aims to replace a heterogeneous material with a homogeneous one that exhibits an equivalent macroscopic behaviour (Engquist2008). For most computational homogenization problems, physics-informed neural networks (PINNs) (RAISSI2019686) currently cannot match the computational efficiency and reliability of classical Finite Element Method (FEM) and Fast Fourier Transform (FFT) approaches (GEERS20102175; Davoli), despite promising theoretical advantages (Grossmann). However, PINNs demonstrate clear advantages for mesh-free computation of complex microstructures, multi-query parametric studies, and scenarios involving irregular geometries where traditional meshing becomes problematic (RODRIGUES201873).
@@
Recent advances in PINN architectures have begun addressing fundamental limitations in multiscale problems (Park2022; Leung2022; Soyarslan2024). These developments, combined with the inherent advantages of mesh-free computation and rapid evaluation for trained networks, position PINNs as valuable tools for specific applications while classical methods remain optimal for computational homogenization workflows in production.
@@
A critical gap remains in establishing rigorous error bounds and robust formulations for elliptic partial differential equations (PDEs) within the PINN framework. While existing PINN approaches primarily rely on the strong form of the PDE residual, this formulation can be sensitive to solution regularity and may not provide reliable error estimates (krishnapriyan2021; deryck2023; mishra2023). Variational PINNs (VPINNs) address some of these limitations by incorporating weak formulations that utilize test functions and integrate the PDE residual over the domain, thereby imposing weaker regularity requirements and potentially enhancing solution stability. However, VPINNs remain sensitive to the choice of test functions and still lack robustness in error estimation (E2018; kharazmi2019).
@@
To address these challenges in PINN-based homogenization, we propose a scheme that enables the calculation of guaranteed upper and lower bounds for effective parameters. Using two microstructure examples, we demonstrate how classic PINNs can fail when applied to piecewise constant materials. Although sophisticated techniques exist for handling discontinuous materials, including GFEM (Legrain), X-FEM (Belytschko_2009), the immersed boundary method (LeVeque), domain decomposition PINNs (AmeyaDJagtap2020), and various adaptive PINN approaches (Liu2023; ZHAO2025118184), robust application of advanced PINNs in homogenization settings remains under development.
@@
Our approach employs both primal and dual formulations of the elliptic PDE governing the homogenization problem, rather than relying solely on the primal formulation. The dual formulation provides complementary information that enables more accurate and reliable bounds on effective material properties, as demonstrated for classical homogenization methods by Gaynutdinova2022; Gaynutdinova2023. For the network architecture, we adopt the frameworks of JIANG2023115972 and Soyarslan2024, which incorporate periodic boundary conditions directly without loss penalty terms. We extend their architecture to solve the dual formulation problem while satisfying the divergence-free constraint (richterpowell2022; farkane2023), then use the primal and dual solutions to derive error bounds. Recognizing that the strong form can still yield poor performance in classic PINNs, we also implement the primal-dual approach for VPINNs and compare results across different test function choices (Berrone2022; ROJAS2024116904). Our results reveal that strong-form PINNs outperform both VPINNs and FEM benchmarks only when materials exhibit sufficiently gradual phase transitions, while VPINNs show sensitivity to test function selection. In all cases, however, the guaranteed bounds to the effective parameters can be utilized to assess the quality of the computed results.
@@
The rest of this work is structured as follows. In Section 2, we recall the zeroth-order homogenization theory for both primal and dual formulations in the steady-state linear 2D heat equation. In Section 3, we introduce the PINN architecture and construct the strong training loss for the dual formulation, as well as the variational loss for both primal and dual formulations. In Subsection 4.1.1, we compare the performance of the classic PINNs for different material functions and address a common mode of PINN failure, whereas in Subsection 4.1.2 we compare the performance of VPINNs for different types and cardinalities of test bases; in both cases, for a single inclusion. Subsection 4.2 then extends this discussion to a microstructure with multiple inclusions. Section 5 summarizes our findings and discusses the pros and cons of each method, and Section 6 provides the outlook for future applications.
@@ 2 Primal and Dual Formulation of Homogenization
In this part, we employ computational homogenization, which determines the effective properties of heterogeneous materials by solving boundary value problems on their representative periodic cells. First, we introduce the governing equation for the homogenization of linear steady-state heat transport in 2D by recalling the strong form of the equation, from which the typical PINN loss is constructed (Park2022; Leung2022; JIANG2023115972; WU2023112521; Soyarslan2024), and introduce its dual form. Then we present the variational forms for both the primal and dual formulations, which are used in most classical numerical methods, and will serve as the basis for the VPINN approach. Finally, we show how to calculate the estimates of the homogenized conductivity matrix, as well as its guaranteed upper and lower bounds.
@@ 2.1 Strong form
Let us consider a periodic unit cell $X\subset\mathbb{R}^{2}$ representing a periodic heterogeneous solid, where its microstructural thermal conductivity second-order tensor can be represented in the matrix notation as a $2\times 2$ symmetric matrix $\bm{A}$. Its spatial variability within the periodic cell is then expressed as a function of the spatial coordinate $\bm{x}$ as $\bm{A}(\bm{x}):X\to\mathbb{R}_{\rm{sym}}^{2\times 2}$, which is uniformly positive definite, with each component essentially bounded in $X$. According to the zeroth-order homogenization theory (MICHEL1999109; jikov2012homogenization), the temperature field $u(\bm{x})$ within the unit cell takes the form
$u(\bm{x})=\bm{\xi}^{\rm{T}}\bm{x}+\widetilde{u}(\bm{x}),\;\bm{x}\in X,$ (1)
where $\bm{\xi}^{\rm{T}}\bm{x}$ represents the averaged contribution for the prescribed macroscopic temperature gradient $\bm{\xi}$ and $\widetilde{u}(\bm{x})$ is the micro-fluctuation part caused by the material heterogeneity. The temperature fluctuation $\widetilde{u}(\bm{x})$ is a $X$-periodic function in the $x_{1}$$x_{2}$ plane, $\bm{x}=[x_{1},x_{2}]$.
Under steady-state conditions with no heat sources, the heat flux $\bm{j}(\bm{x})$ is related to the temperature gradient $\bm{e}(\bm{x})=\nabla u(\bm{x})=[\frac{\partial u}{\partial x_{1}},\frac{\partial u}{\partial x_{2}}]^{\rm T}$ by Fourier’s law as:
$\bm{j}(\bm{x})=-\bm{A}(\bm{x})\bm{e}(\bm{x})=-\bm{A}(\bm{x})\nabla u(\bm{x}),$ (2)
and the divergence of the heat flux must be zero, i.e.,
$\nabla\cdot\bm{j}(\bm{x})=0.$ (3)
The strong form of the governing differential equation of the steady-state heat conduction then requires one to find $\widetilde{u}\in C^{2}_{\rm per}(X)$ such that:
@@
where $\nabla\cdot$ denotes the divergence operator, i.e., $\nabla\cdot\bm{f}=\frac{\partial f_{1}}{\partial x_{1}}+\frac{\partial f_{2}}{\partial x_{2}}$ and $C^{2}_{\rm per}(X)$ denotes the space of twice-differentiable periodic continuous function on $X$.
The dual form of Eq. (4) is defined for the thermal resistivity $\bm{A}^{-1}(\bm{x})$ (VONDREJC2015). For the heat transfer problem, we can derive it analogously from the property of the temperature gradient field $\bm{e}$ (briane2008duality; milton2022theory), following the Helmholtz decomposition theorem (jikov2012homogenization), i.e.:
$\nabla\times\bm{e}(\bm{x})=\bm{0},\;\forall\bm{x}\in X.$ (5)
@@
$\widetilde{\bm{v}}(\bm{x})=\bm{Q}\nabla\widetilde{w}(\bm{x}),$ (9)
where $\bm{Q}$ is a $90^{\circ}$ rotation matrix in the $x_{1}$$x_{2}$ plane (Girault1979; milton2022theory) of the form
$\bm{Q}=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$. (10)
The divergence of vector $\widetilde{\bm{v}}(\bm{x})$ can then be written as
$\nabla\cdot\widetilde{\bm{v}}(\bm{x})=\left[\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{2}}\right]\cdot\begin{bmatrix}-\frac{\partial\widetilde{w}}{\partial x_{2}}\\ \frac{\partial\widetilde{w}}{\partial x_{1}}\end{bmatrix}=-\frac{\partial^{2}\widetilde{w}}{\partial x_{2}\partial x_{1}}+\frac{\partial^{2}\widetilde{w}}{\partial x_{1}\partial x_{2}},$ (11)
which is always equal to zero, thanks to the symmetry of partial derivatives. Then, the problem (7) can be written as to find $\widetilde{w}(\bm{x})\in C^{2}_{\rm per}(X)$ such that
@@
with the divergence-free constraint (8) is automatically fulfilled. Note that in order to apply the curl operator $\nabla\times$ to a 2D vector function $\bm{f}$, the function has to be expanded to a 3D vector function $f_{1},f_{2},f_{3}=0]$. The condition $\nabla\times\bm{f}=\bm{0}$ can then be written as
$\left[\frac{\partial f_{3}}{\partial x_{2}}-\frac{\partial f_{2}}{\partial x_{3}},\frac{\partial f_{1}}{\partial x_{3}}-\frac{\partial f_{3}}{\partial x_{1}},\frac{\partial f_{2}}{\partial x_{1}}-\frac{\partial f_{1}}{\partial x_{2}}\right]=\bm{0}.$ (13)
Since $f_{3}=0$ and $\bm{f}$ is a function in the $x_{1}$$x_{2}$ plane, the first two components are always zero, and the equation (12) simplifies to contain only the last component, thus reducing to a scalar problem similar to Eq. (4). We keep the original notation for brevity, but from now on, when we write $\nabla\times\bm{f}$, we actually mean its scalar form $\frac{\partial f_{2}}{\partial x_{1}}-\frac{\partial f_{1}}{\partial x_{2}}$.
@@ 2.2 Weak form
Despite promising theoretical advantages, strong-form PINNs currently cannot match the computational efficiency and reliability of classical computational approaches for most PDEs (Grossmann). As the analytical solution of problems (4) and (12) is usually unavailable, most frameworks, such as the finite element method (FEM) (yvonnet2019computational) and Fourier-Galerkin method (moulinec1994fast; schneider2021review), rely on their respective variational (weak) forms. This allows for the relaxation of the solution regularity requirement from $C^{2}_{\rm per}(X)$ to space $H^{1}_{\rm per}(X)=\{\bm{u}\in L^{2}(X,\mathbb{R});\bm{u}_{\rm per}\in H^{1}_{\rm loc}(\mathbb{R}^{2},\mathbb{R})\}$, where $\bm{u}_{\rm per}$ is the $X$-periodic extension of $\bm{u}$ and $H^{1}_{\rm loc}(\mathbb{R}^{2},\mathbb{R})$ are functions with squared gradients integrable on every compact subset of $\mathbb{R}^{2}$. Instead of solving (4), we want to find $\widetilde{u}\in H^{1}_{\rm per}(X)$ such that
$\int_{X}(\nabla\phi(\bm{x}))^{\rm{T}}\bm{A}(\bm{x})(\bm{\xi}+\nabla\widetilde{u}(\bm{x})){\rm d}\bm{x}=0,\;\;\text{for all }\phi\in H^{1}_{\rm per}(X),$ (14)
where $\phi(\bm{x})$ is a test function. The boundary term $\int_{\partial X}(n\cdot\nabla\phi)\widetilde{u}(\bm{x}){\rm d}s=0$, which appears alongside (14) after the integration by parts, vanishes automatically, since it has opposite values on the opposite boundaries of $X$ due to periodicity. Similarly, the variational form of (12) is to find $\widetilde{w}\in H^{1}_{\rm per}(X)$ such that
$\int_{X}(\bm{Q}\nabla\psi(\bm{x}))^{\rm{T}}\;\bm{A}^{-1}(\bm{x})(\bm{\xi}+\bm{Q}\nabla\widetilde{w}(\bm{x})){\rm d}\bm{x}=0,\;\;\text{for all }\psi\in H^{1}_{\rm per}(X),$ (15)
with a test function $\psi(\bm{x})$$\in H^{1}_{\rm per}(X)$.
@@ 2.3 Upper and lower bound to homogenized parameters
The exact solution of (14) is also the minimizer of the related homogenization problem (jikov2012homogenization), which reads to find a homogenized coefficient matrix $\bm{A}^{*}\in{\mathbb{R}}^{2\times 2}$ such that
@@
Conversely, the inverse $\bm{B}^{*}$ of the homogenized coefficient $\bm{A}^{*}$ can be calculated using the dual formulation (haslinger1995optimum; VONDREJC2015; Gaynutdinova2022):
@@
Given the variational structure of (16) and (17), for which the true minimizer exists in the infinite-dimensional space $H^{1}_{\rm per}(X)$, substituting any minimizer over subspaces of $H^{1}_{\rm per}(X)$ (i.e., approximate solutions $\widetilde{u}_{h}$ and $\widetilde{w}_{h}$) into the integrals in (16) and (17), respectively, for given $\bm{\xi}$ and $\bm{\zeta}$ yields larger values of the effective coefficients than for the true minimizers, and thus it yields the upper bounds $\bm{A}^{*}_{h}$ and $\bm{B}^{*}_{h}$ to $\bm{A}^{*}$ and $\bm{B}^{*}$ in the sense that
@@
In practice, we solve (4) or (14) (as well as (12) and (15)) for given $\bm{\xi}$ and $\bm{\zeta}\in\{[1,0]^{\rm T},[0,1]^{\rm T}\}$ to compute the diagonal elements of $\bm{A}^{*}_{h}$ and $\bm{B}^{*}_{h}$.